# What’s in a Name? Auditing Large Language Models for Race and Gender Bias

### v2

@@ Appendix F Descriptive Statistics
Scenario Variation Context Level Mean Black White Male Female Black Men White Men Black Women White Women Sports Basketball Low 55 53 55 53 56 54 54 52 [55, 55] [53, 54] [55, 55] [53, 53] [56, 56] [54, 55] [54, 55] [52, 52] High 85 79 82 82 86 79 85 79 [85, 85] [78, 79] [82, 83] [81, 82] [85, 86] [78, 80] [84, 85] [78, 79] Numeric 56 56 56 56 56 56 56 56 [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] Football Low 57 56 58 56 58 58 57 55 [57, 58] [56, 56] [58, 58] [56, 56] [58, 58] [58, 58] [57, 57] [54, 55] High 61 60 61 61 62 60 61 60 [61, 62] [60, 60] [60, 61] [60, 61] [61, 62] [60, 60] [61, 62] [59, 60] Numeric 56 56 56 56 56 56 56 56 [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] Hockey Low 55 56 57 54 56 57 54 54 [55, 55] [55, 56] [56, 57] [54, 54] [56, 56] [57, 57] [53, 54] [54, 54] High 81 79 80 80 81 79 81 79 [80, 81] [78, 79] [79, 80] [79, 81] [80, 81] [78, 79] [81, 82] [78, 80] Numeric 56 56 56 56 56 56 56 56 [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] Lacrosse Low 56 54 56 54 56 55 55 54 [55, 56] [54, 55] [56, 56] [54, 54] [56, 57] [55, 56] [54, 55] [53, 54] High 70 70 69 71 69 69 72 71 [70, 71] [70, 71] [69, 70] [71, 72] [68, 70] [68, 70] [71, 72] [70, 72] Numeric 56 56 56 56 56 56 56 56 [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] Note: This table displays the mean and confidence intervals (enclosed in brackets) for all the responses collected in the Sports scenario. It provides descriptive statistics to compare across races and genders, incorporating intersectional analysis.
Table 8: Sports

### v3

@@ Appendix F Descriptive Statistics
Table 8: Sports Scenario Variation Context Level Mean Black White Male Female Black Men White Men Black Women White Women Sports Basketball Low 55 53 55 53 56 54 54 52 [55, 55] [53, 54] [55, 55] [53, 53] [56, 56] [54, 55] [54, 55] [52, 52] High 85 79 82 82 86 79 85 79 [85, 85] [78, 79] [82, 83] [81, 82] [85, 86] [78, 80] [84, 85] [78, 79] Numeric 56 56 56 56 56 56 56 56 [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] Football Low 57 56 58 56 58 58 57 55 [57, 58] [56, 56] [58, 58] [56, 56] [58, 58] [58, 58] [57, 57] [54, 55] High 61 60 61 61 62 60 61 60 [61, 62] [60, 60] [60, 61] [60, 61] [61, 62] [60, 60] [61, 62] [59, 60] Numeric 56 56 56 56 56 56 56 56 [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] Hockey Low 55 56 57 54 56 57 54 54 [55, 55] [55, 56] [56, 57] [54, 54] [56, 56] [57, 57] [53, 54] [54, 54] High 81 79 80 80 81 79 81 79 [80, 81] [78, 79] [79, 80] [79, 81] [80, 81] [78, 79] [81, 82] [78, 80] Numeric 56 56 56 56 56 56 56 56 [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] Lacrosse Low 56 54 56 54 56 55 55 54 [55, 56] [54, 55] [56, 56] [54, 54] [56, 57] [55, 56] [54, 55] [53, 54] High 70 70 69 71 69 69 72 71 [70, 71] [70, 71] [69, 70] [71, 72] [68, 70] [68, 70] [71, 72] [70, 72] Numeric 56 56 56 56 56 56 56 56 [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] [56, 56] Note: This table displays the mean and confidence intervals (enclosed in brackets) for all the responses collected in the Sports scenario. It provides descriptive statistics to compare across racial and gender groups.